# arXiv:2603.02631v1[cs.CL] 03 Mar 2026

### v1

@@ 1 Introduction
Prompt compression therefore, is an attractive avenue to reduce LLM inference cost. Recent work on Speculative Prefill Liu et al. (2025) addresses this problem by estimating token-level importance scores using a lightweight draft model. In this setting, the target model is the large language model used to perform the final task inference on the compressed input prompt while the draft model is a smaller model used at inference time to estimate token importance to guide prompt compression. Both models belong to the same model family. In practice however, some frontier models don’t have smaller in-family draft models available (eg: DeepSeek-V3.1/R1 (5; D. Guo, D. Yang, H. Zhang, J. Song, P. Wang, Q. Zhu, R. Xu, R. Zhang, S. Ma, X. Bi, X. Zhang, X. Yu, Y. Wu, Z. F. Wu, Z. Gou, Z. Shao, Z. Li, Z. Gao, A. Liu, B. Xue, B. Wang, B. Wu, B. Feng, C. Lu, C. Zhao, C. Deng, C. Ruan, D. Dai, D. Chen, D. Ji, E. Li, F. Lin, F. Dai, F. Luo, G. Hao, G. Chen, G. Li, H. Zhang, H. Xu, H. Ding, H. Gao, H. Qu, H. Li, J. Guo, J. Li, J. Chen, J. Yuan, J. Tu, J. Qiu, J. Li, J. L. Cai, J. Ni, J. Liang, J. Chen, K. Dong, K. Hu, K. You, K. Gao, K. Guan, K. Huang, K. Yu, L. Wang, L. Zhang, L. Zhao, L. Wang, L. Zhang, L. Xu, L. Xia, M. Zhang, M. Zhang, M. Tang, M. Zhou, M. Li, M. Wang, M. Li, N. Tian, P. Huang, P. Zhang, Q. Wang, Q. Chen, Q. Du, R. Ge, R. Zhang, R. Pan, R. Wang, R. J. Chen, R. L. Jin, R. Chen, S. Lu, S. Zhou, S. Chen, S. Ye, S. Wang, S. Yu, S. Zhou, S. Pan, S. S. Li, S. Zhou, S. Wu, T. Yun, T. Pei, T. Sun, T. Wang, W. Zeng, W. Liu, W. Liang, W. Gao, W. Yu, W. Zhang, W. L. Xiao, W. An, X. Liu, X. Wang, X. Chen, X. Nie, X. Cheng, X. Liu, X. Xie, X. Liu, X. Yang, X. Li, X. Su, X. Lin, X. Q. Li, X. Jin, X. Shen, X. Chen, X. Sun, X. Wang, X. Song, X. Zhou, X. Wang, X. Shan, Y. K. Li, Y. Q. Wang, Y. X. Wei, Y. Zhang, Y. Xu, Y. Li, Y. Zhao, Y. Sun, Y. Wang, Y. Yu, Y. Zhang, Y. Shi, Y. Xiong, Y. He, Y. Piao, Y. Wang, Y. Tan, Y. Ma, Y. Liu, Y. Guo, Y. Ou, Y. Wang, Y. Gong, Y. Zou, Y. He, Y. Xiong, Y. Luo, Y. You, Y. Liu, Y. Zhou, Y. X. Zhu, Y. Huang, Y. Li, Y. Zheng, Y. Zhu, Y. Ma, Y. Tang, Y. Zha, Y. Yan, Z. Z. Ren, Z. Ren, Z. Sha, Z. Fu, Z. Xu, Z. Xie, Z. Zhang, Z. Hao, Z. Ma, Z. Yan, Z. Wu, Z. Gu, Z. Zhu, Z. Liu, Z. Li, Z. Xie, Z. Song, Z. Pan, Z. Huang, Z. Xu, Z. Zhang, and Z. Zhang (2025)), Kimi-K2 Team et al. (2025) etc.) Moreover, agentic systems frequently combine heterogeneous models due to cost, availability, and deployment constraints. This raises a natural question: does attention-based token importance generalize across model families?
@@
In this work, we show that attention-based speculative prefill generalizes across model families, and it can be used as a reliable prompt compression mechanism for target models which don’t have a supporting in-family draft model. Importantly, we do not modify the original speculative prefill algorithm itself: we use the same attention-based token scoring and chunk selection as prior work Liu et al. (2025). The only difference is that the draft and target models are drawn from different model families, including combinations of Qwen, LLaMA, and DeepSeek models (Yang et al., 2025; Grattafiori et al., 2024; Guo et al., 2025; Liu et al., 2024).
@@
  • We evaluate our method on long context benchmarks like LongBench Bai et al. (2024; 2025), RULER Hsieh et al. (2024) along with some code debugging tasks Zhang et al. (2024b). Across a wide range of tasks, cross-model prompt compression largely retains 90–100% of full-prompt baseline performance, despite substantial differences in model architecture and tokenizers between draft and target models.
@@ 5.1 Cross-Family Speculative Prefill Preserves Task Performance
Results for LongBench v2 and Code Debug are presented in Tables 2 and 2, with additional results on LongBench v1 and RULER reported in Appendix A.4. We use greedy decoding for all reported results. For Code Debug, the average input length for DeepSeek-R1 is $\sim$110k tokens, and keep rates of 30%, 20%, and 15% reduce the effective context to roughly 33k, 22k, and 16k tokens respectively. For LongBench v2, the average input length is 242k, 260k and 248k tokens under LLaMA-8B, Qwen3-8B and DeepSeek-R1 as target models, and the compressed lengths follow the keep rates specified in Table 2.
@@
Across diverse benchmarks, cross-model speculative prefill maintains performance close to the full-prompt baseline (typically within 90–100%), even at low keep rates of 6% on selected Longbench V2 tasks. This robustness holds despite differences in architecture and tokenization between the draft and target model families, suggesting that token importance estimates are not strongly tied to a specific model family. We see some degradation on Code Debug under aggressive compression (e.g., 15% keep rate), which is discussed in Appendix A.5. In some cases, most notably on LongBench v2, compressed prompts achieve performance slightly exceeding the full-prompt baseline (Table  2). We attribute this to the denoising effect of context compression: by removing large amounts of irrelevant context, speculative prefill simplifies the problem.
@@
For RULER with input context lengths of 128k (Table  4), we note that the full-prompt baseline is evaluated using SnapStream KV cache compression Li et al. (2025) on our custom Reconfigurable Dataflow Unit (RDU) hardware. This is needed to make 128k-token inference feasible within on-chip memory constraints for DeepSeek-V3. This cache compression contributes to loss of long range information needed for retrieval-heavy tasks in RULER, thus observing the lower baseline performance. In contrast, speculative prefill reduces the effective context length prior to inference, thus proactively removing noisy context and mitigating this degradation. Together, these results highlight that cross model speculative prefill can preserve accuracy and in some cases, improve task performance under practical deployment constraints. Overall, cross-model speculative prefill enables substantial prompt compression while largely preserving downstream task performance.
@@ 5.2 Cross model prompt compression reduces long context cost
5.3 Robustness to Draft Model Selection
We next evaluate the robustness of prompt compression to the choice of draft model while holding the target model fixed. Results on LongBench v2 (Table 2) indicate that effective compression can be achieved using draft models with substantially different parameter scales. For instance, with LLaMA-8B as the target model, both Qwen3-1.7B and Qwen3-0.6B produce compressed prompts that preserve performance close to the full-prompt baseline. Likewise for DeepSeek-R1, competitive results are obtained using either Qwen3-4B or LLaMA-8B as draft models. Overall, performance varies only modestly across draft model choices, suggesting that the compression quality is not strongly dependent on a specific architecture or model scale beyond a minimal effectiveness threshold. This robustness enables flexible draft model selection based on computational cost or resource constraints without sacrificing downstream accuracy.
5.4 Decoupling Target Context Length from Deployment Limits
6 Conclusion
Appendix A Appendix
A.1 Token Importance Estimation and Chunk Selection
A.1.1 Attention-Based Token Importance Estimation
Given an input prompt of length S, the draft model processes the full context and computes attention scores using a fixed number of lookahead tokens. As in prior work Liu et al. (2025), lookahead is used to mitigate positional biases in attention distributions by aggregating signals from multiple decoding steps. Let the attention tensor be of shape [N, L, S, H], where N is the number of lookahead tokens, L the number of transformer layers, S the prompt length, and H the number of attention heads. We compute token importance scores using the same max–mean aggregation strategy as speculative prefill: attention scores are first maximized over the layer and head dimensions, and then averaged across the lookahead dimension. This produces a single scalar importance score for each input token.
A.1.2 Chunk-Based Prompt Compression
A.2 Benchmark details
LongBench v1 and v2:
RULER and InfiniteBench:
A.3 Implementation Details
A.4 Additional Results
We report additional results on LongBench v1 and RULER in Tables 3 and 4. On LongBench v1, we apply cross-model speculative prefill to compress inputs to a fixed target length of 4k tokens for each target model, corresponding to an average keep rate of 45%. Across the selected LongBench v1 datasets, the average input length is 8805 tokens for LLaMA-8B and 8867 tokens for Qwen3-8B as target models. Draft models from both the Qwen and LLaMA families act as effective compressors, with cross-family prompt compression achieving accuracies close to the full-prompt baseline.
For RULER, we compress 128k-token inputs down to 16k tokens (12.5% keep rate). Notably, the full-prompt baseline underperforms speculative prefill in this setting. We attribute this to two main factors. First, speculative prefill performs implicit context denoising by retaining only salient regions of the input, which improves accuracy on retrieval-heavy, needle-in-a-haystack tasks. This denoising is missing for the reported baseline results. Second, making 128k full-prompt inference feasible with DeepSeek V3 requires combining it with SnapStream KV cache compression Li et al. (2025) in our deployment setting. While necessary for memory reasons, KV cache compression can discard long-range information, leading to degradation in the full-prompt baseline, particularly for RULER which has retrieval heavy tasks.

### v2

# Cross-Family Speculative Prefill: Training-Free Long-Context Compression with Small Draft Models

@@ 1 Introduction
Prompt compression therefore, is an attractive avenue to reduce LLM inference cost. Recent work on Speculative Prefill (Liu et al., 2025) addresses this problem by estimating token-level importance scores using a lightweight draft model. In this setting, the target model is the large language model used to perform the final task inference on the compressed input prompt while the draft model is a smaller model used at inference time to estimate token importance to guide prompt compression. Both models belong to the same model family. In practice however, some frontier models don’t have smaller in-family draft models available (eg: DeepSeek-V3.1/R1, Kimi-K2 etc (5; D. Guo, D. Yang, H. Zhang, J. Song, P. Wang, Q. Zhu, R. Xu, R. Zhang, S. Ma, X. Bi, X. Zhang, X. Yu, Y. Wu, Z. F. Wu, Z. Gou, Z. Shao, Z. Li, Z. Gao, A. Liu, B. Xue, B. Wang, B. Wu, B. Feng, C. Lu, C. Zhao, C. Deng, C. Ruan, D. Dai, D. Chen, D. Ji, E. Li, F. Lin, F. Dai, F. Luo, G. Hao, G. Chen, G. Li, H. Zhang, H. Xu, H. Ding, H. Gao, H. Qu, H. Li, J. Guo, J. Li, J. Chen, J. Yuan, J. Tu, J. Qiu, J. Li, J. L. Cai, J. Ni, J. Liang, J. Chen, K. Dong, K. Hu, K. You, K. Gao, K. Guan, K. Huang, K. Yu, L. Wang, L. Zhang, L. Zhao, L. Wang, L. Zhang, L. Xu, L. Xia, M. Zhang, M. Zhang, M. Tang, M. Zhou, M. Li, M. Wang, M. Li, N. Tian, P. Huang, P. Zhang, Q. Wang, Q. Chen, Q. Du, R. Ge, R. Zhang, R. Pan, R. Wang, R. J. Chen, R. L. Jin, R. Chen, S. Lu, S. Zhou, S. Chen, S. Ye, S. Wang, S. Yu, S. Zhou, S. Pan, S. S. Li, S. Zhou, S. Wu, T. Yun, T. Pei, T. Sun, T. Wang, W. Zeng, W. Liu, W. Liang, W. Gao, W. Yu, W. Zhang, W. L. Xiao, W. An, X. Liu, X. Wang, X. Chen, X. Nie, X. Cheng, X. Liu, X. Xie, X. Liu, X. Yang, X. Li, X. Su, X. Lin, X. Q. Li, X. Jin, X. Shen, X. Chen, X. Sun, X. Wang, X. Song, X. Zhou, X. Wang, X. Shan, Y. K. Li, Y. Q. Wang, Y. X. Wei, Y. Zhang, Y. Xu, Y. Li, Y. Zhao, Y. Sun, Y. Wang, Y. Yu, Y. Zhang, Y. Shi, Y. Xiong, Y. He, Y. Piao, Y. Wang, Y. Tan, Y. Ma, Y. Liu, Y. Guo, Y. Ou, Y. Wang, Y. Gong, Y. Zou, Y. He, Y. Xiong, Y. Luo, Y. You, Y. Liu, Y. Zhou, Y. X. Zhu, Y. Huang, Y. Li, Y. Zheng, Y. Zhu, Y. Ma, Y. Tang, Y. Zha, Y. Yan, Z. Z. Ren, Z. Ren, Z. Sha, Z. Fu, Z. Xu, Z. Xie, Z. Zhang, Z. Hao, Z. Ma, Z. Yan, Z. Wu, Z. Gu, Z. Zhu, Z. Liu, Z. Li, Z. Xie, Z. Song, Z. Pan, Z. Huang, Z. Xu, Z. Zhang, and Z. Zhang (2025); K. Team, Y. Bai, Y. Bao, G. Chen, J. Chen, N. Chen, R. Chen, Y. Chen, Y. Chen, Y. Chen, et al. (2025))). Moreover, agentic systems frequently combine heterogeneous models due to cost, availability, and deployment constraints. This raises a natural question: does attention-based token importance generalize across model families?
@@
In this work, we show that attention-based speculative prefill generalizes across model families, and it can be used as a reliable prompt compression mechanism for target models which don’t have a supporting in-family draft model. Importantly, we do not modify the original speculative prefill algorithm itself: we use the same attention-based token scoring and chunk selection as prior work (Liu et al., 2025). The only difference is that the draft and target models are drawn from different model families, including combinations of Qwen, LLaMA, and DeepSeek models (Yang et al., 2025; Grattafiori et al., 2024; Guo et al., 2025; Liu et al., 2024).
@@
  • We evaluate our method on long context benchmarks like LongBench, RULER along with some code debugging tasks. Across a wide range of tasks, cross-model prompt compression largely retains 90–100% of full-prompt baseline performance, despite substantial differences in model architecture and tokenizers between draft and target models(Bai et al., 2024; 2025; Hsieh et al., 2024; Zhang et al., 2024b).
@@ 3.1 Keep Rate Selection
For most of our results reported in Section 5, we select keep rates that correspond to compressed prompt lengths of 8k, 16k, 24k, and 32k tokens, which are compatible with the paged-attention execution scheme on our custom hardware Reconfigurable Dataflow Unit (RDU). In particular, KV caches are stored in 4k-token blocks, and attention computation operates over these blocks during inference. Choosing compressed prompt lengths that are multiples of this block size ensures efficient KV cache paging and streaming during attention computation.
@@ 5.1 Cross-Family Speculative Prefill Preserves Task Performance
Results for LongBench v2 and Code Debug are presented in Tables 2 and 2, with additional results on LongBench v1 and RULER reported in Tables 3 and 4 respectively. We use greedy decoding for all reported results. For Code Debug, the average input length for DeepSeek-R1 is $\sim$110k tokens, and keep rates of 30%, 20%, and 15% reduce the effective context to roughly 33k, 22k, and 16k tokens respectively. For LongBench v2, the average input length is 242k, 260k and 248k tokens under LLaMA-8B, Qwen3-8B and DeepSeek-R1 as target models, and the compressed lengths follow the keep rates specified in Table 2.
@@
Across diverse benchmarks, cross-model speculative prefill maintains performance close to the full-prompt baseline (typically within 90–100%), even at low keep rates of 6% on selected Longbench V2 tasks. This robustness holds despite differences in architecture and tokenization between the draft and target model families, suggesting that token importance estimates are not strongly tied to a specific model family. For LLaMA-8B, we evaluate both an in-family draft (LLaMA-1B) and a cross-family draft (Qwen3-1.7B). The results show comparable performance across keep rates (e.g., 29.4% vs 29.6% at 25% keep rate, and 30.6% vs 31.2% at 50%). Similar trends are observed for Qwen3-8B when paired with LLaMA-1B and Qwen3-1.7B drafts. These results indicate that speculative prefill is not strongly dependent on cross-family pairing, and that small draft models of comparable size provide similar token saliency signals for prompt compression.
For DeepSeek-R1, we additionally explore keep rates of 3%, 6%, and 10% on LongBench-v2. These correspond to compressed prompt lengths of approximately 8k, 16k, and 24k tokens, respectively. We observe that performance improves as the keep rate increases from 3% to 6%, indicating the expected trade-off between compression and task performance. The accuracy decreases slightly at 10%, suggesting a diminishing return from retaining additional context. Finally, we note that the larger draft model (Qwen3-4B-Instruct-2507) achieves higher performance on DeepSeek-R1 because it supports a 262k native context length, compared to 32k for Qwen3-1.7B. The longer context allows the draft model to estimate attention over long inputs more reliably, which improves token selection during compression.
We see some degradation on Code Debug under aggressive compression (e.g., 15% keep rate), which is discussed in Appendix A.5. In some cases, most notably on LongBench v2, compressed prompts achieve performance slightly exceeding the full-prompt baseline (Table  2). We attribute this to the denoising effect of context compression: by removing large amounts of irrelevant context, speculative prefill simplifies the problem.
@@
For RULER with input context lengths of 128k (Table  4), we note that the full-prompt baseline is evaluated using SnapStream KV cache compression (Li et al., 2025) on our custom RDU hardware. This is needed to make 128k-token inference feasible within on-chip memory constraints for DeepSeek-V3. This cache compression contributes to loss of long range information needed for retrieval-heavy tasks in RULER, thus observing the lower baseline performance. In contrast, speculative prefill reduces the effective context length prior to inference, thus proactively removing noisy context and mitigating this degradation. Together, these results highlight that cross model speculative prefill can preserve accuracy and in some cases, improve task performance under practical deployment constraints. Overall, cross-model speculative prefill enables substantial prompt compression while largely preserving downstream task performance. Additional details are given in A.4.
@@ 5.2 Cross model prompt compression reduces long context cost
5.3 Robustness to Draft Model Selection
We next evaluate the robustness of prompt compression to the choice of draft model while holding the target model fixed. Results on LongBench v2 (Table 2) indicate that effective compression can be achieved using draft models with substantially different parameter scales. For instance, with LLaMA-8B as the target model, both Qwen3-1.7B and Qwen3-0.6B produce compressed prompts that preserve performance close to the full-prompt baseline. Likewise for DeepSeek-R1, competitive results are obtained using either Qwen3-4B or LLaMA-8B as draft models. Our experiments show that small draft models (1B–1.7B range) already provide sufficiently reliable saliency estimates for 8B target models. Larger draft models (e.g., Qwen3-4B) improve performance primarily when processing very long inputs, where stronger long-context modeling leads to more accurate attention estimation. Overall, performance varies only modestly across draft model choices, suggesting that the compression quality is not strongly dependent on a specific architecture or model scale beyond a minimal effectiveness threshold. Tokens that are important for reasoning (numbers, entities, key instructions) tend to attract attention consistently across different LLM architectures trained on similar data. This robustness enables flexible draft model selection based on computational cost or resource constraints without sacrificing downstream accuracy.
5.4 Decoupling Target Context Length from Deployment Limits
6 Conclusion
Appendix A Appendix
A.1 Token Importance Estimation and Chunk Selection
A.1.1 Attention-Based Token Importance Estimation
Given an input prompt of length S, the draft model processes the full context and computes attention scores using a fixed number of lookahead tokens. As in prior work (Liu et al., 2025), lookahead is used to mitigate positional biases in attention distributions by aggregating signals from multiple decoding steps. Let the attention tensor be of shape [N, L, S, H], where N is the number of lookahead tokens, L the number of transformer layers, S the prompt length, and H the number of attention heads. We compute token importance scores using the same max–mean aggregation strategy as speculative prefill: attention scores are first maximized over the layer and head dimensions, and then averaged across the lookahead dimension. This produces a single scalar importance score for each input token.
A.1.2 Chunk-Based Prompt Compression
A.2 Benchmark details
LongBench v1 and v2:
RULER and InfiniteBench:
A.3 Implementation Details
A.4 Additional result Analysis
We report results on LongBench v1 and RULER in Tables 3 and 4. On LongBench v1, we apply cross-model speculative prefill to compress inputs to a fixed target length of 4k tokens for each target model, corresponding to an average keep rate of 45%. Across the selected LongBench v1 datasets, the average input length is 8805 tokens for LLaMA-8B and 8867 tokens for Qwen3-8B as target models. Draft models from both the Qwen and LLaMA families act as effective compressors, with cross-family prompt compression achieving accuracies close to the full-prompt baseline.
For RULER, we compress 128k-token inputs down to 16k tokens (12.5% keep rate). Notably, the full-prompt baseline underperforms speculative prefill in this setting. We attribute this to two main factors. First, speculative prefill performs implicit context denoising by retaining only salient regions of the input, which improves accuracy on retrieval-heavy, needle-in-a-haystack tasks. This denoising is missing for the reported baseline results. Second, making 128k full-prompt inference feasible with DeepSeek V3 requires combining it with SnapStream KV cache compression (Li et al., 2025) in our deployment setting. While necessary for memory reasons, KV cache compression can discard long-range information, leading to degradation in the full-prompt baseline, particularly for RULER which has retrieval heavy tasks.